the petty-cash drawer of a small office containing $16.25. If there were twice as many nickels as quarters and as many dimes as nickels and quarters combined, how many coins of each type were there?

Answer by mananth(16949)

let quarters be x
nickels =2x
dimes = (2x+x)
..
Total money = 16.25
25x +5*2x+10*(2x+x)=1625
25x+10x+20x+10x=1625
65x=1625
x= 1625/65
x= 25 the number of quarters
nickels = 2x = 2*25=50
dimes = 3x=3*25=75